A projectile is fired at an angle of 60 degrees with initial velocity of 100m/s find, time of asent


1
Expert's answer
2021-03-31T09:38:54-0400

In component form, initial velocity "u_o=100sin60^0"

time of ascent "=\frac{2u_{o}}{g}=\frac{2\times 100sin60^{0}}{9.8}= 17.67s"